# suppSupplementary References

### v2

@@ 4.2 Deterministic inference mode
We evaluate MEGA on the HMR task using the deterministic inference mode defined in Sec. 3.4, and compare its performance to SOTA methods on 3DPW [55] and EMDB [35]. Results for other methods are taken from the corresponding papers when available or computed using official implementations and model weights. To ensure fairness, we only compare methods trained on standard datasets (see Sec. 4.1), with the exception of TokenHMR [19], which uses additional data but is included in Tab. 1 because it also uses mesh tokenization. Note that the results for FastMETRO [11] and Virtual Marker [53] differ from those in their original papers because we present results without finetuning on the 3DPW dataset. Non-parametric methods predicting 3D coordinates often struggle to generalize to unseen datasets, and skipping finetuning on the 3DPW training set significantly degrades their performance.
MEGA outperforms all other methods on in-the-wild datasets. Using an HRNet [79] backbone, MEGA significantly surpasses both parametric [37, 46] and non-parametric [11, 53, 21] methods, especially on the EMDB [35] dataset. With a ViT [17] backbone, MEGA also achieves SOTA performance. However, comparisons with TokenHMR [19] are not completely fair, as this method is trained with additional 2D data and Bedlam [5]. We also provide results of MEGA without pre-training on motion capture data, demonstrating the interest of the self-supervised pre-training step.
@@ 4.3 Stochastic inference mode
We evaluate MEGA in stochastic inference mode (see Sec. 3.4) for the HMR task on the 3DPW [55] dataset, comparing our performance with SOTA multi-output approaches. Results from other methods are obtained from their respective papers. We exclusively compare methods trained on standard datasets (see Sec. 4.1), providing a version of MEGA using a ResNet-50 [28] backbone for fair comparisons. Following prior works [10, 4, 41], we assess accuracy using standard metrics computed with the minimum error sample out of $Q$ predictions, with $Q$ ranging from $1$ to $25$. We also calculate the relative improvement between 1 and 25 samples. Results are presented in Tab. 2.
MEGA achieves SOTA performance across all metrics and sample sizes. Notably, MEGA exhibits significantly higher accuracy than other methods when generating a single sample, and consistently demonstrates the best or second-best relative improvement among methods using a ResNet-50 backbone. With an HRNet [79] backbone, MEGA outperforms all probabilistic methods and surpasses SOTA single-output methods, even with a single stochastic generation. It is worth noting that it takes between 10 and 25 stochastic samples to outperform the deterministic generation (MEGA det), highlighting the utility of deterministic mode for quick and accurate predictions and the advantage of multiple predictions for enhanced accuracy. For a detailed comparison between deterministic and stochastic modes, please refer to Appendix B.
4.4 Random meshes generation
4.5 Stochastic inference mode analysis
We evaluate the correlation between the PVE in deterministic mode and the uncertainty of the predictions in stochastic mode, quantified by the standard deviation of the predicted vertices. We conduct this experiment on 3DPW-OCC [84], which features numerous occlusions, thereby increasing prediction uncertainty. Fig. 5 features multiple samples for images with occlusions, accompanied by a quantitative evaluation in deterministic mode on 3DPW-OCC in Appendix A. In Fig. 4, we visually represent the PVE and the uncertainty as colormaps on meshes. Interestingly, we observe a striking similarity between them, suggesting that the estimated uncertainty could be used to evaluate the model’s per-vertex performance.
Figure 4: PVE and uncertainty. We visualize the PVE and the standard deviation in colormaps on meshes.
Figure 5: Multiple predictions given an image. We generate multiple human meshes given images with occlusions [84] using MEGA in stochastic mode. Note that, while not explicitly shown in the figure, all of these meshes practically have identical image projections.
@@ 5 Conclusion
In this work, we explored self-supervised learning and masked generative modeling on human meshes for HMR. We introduced MEGA, a masked generative autoencoder designed to generate human meshes as discrete token sequences. MEGA’s flexible architecture and generation scheme enable the generation of diverse and realistic meshes and support both single and multi-output HMR. A thorough evaluation demonstrates significant improvement over SOTA in both these domains.
Limitations. While MEGA generally produces accurate predictions, it struggles with extreme poses significantly divergent from the training data. Although uncertainty estimation shows promise, further analysis is required before integrating them into model training supervision. Fig. 8 in the Appendix provides visualizations of these failure cases. Additionally, the experiments were conducted only once, without testing different seeds to confirm the results’ significance. However, it allowed us to significantly reduce our research’s environmental impact.
Future work. MEGA’s adaptable framework suggests potential applications beyond its current scope. Future research could explore generating human meshes conditioned on text inputs [15]. We could also complement image embedding with more observations, such as 2D pose tokens [23] or tokenized meshes of other individuals to model social interactions [26, 58]. Extending this work to videos by incorporating temporal masking during training [64, 65] or including more extreme poses in training data [73] may further improve performance. Additionally, optimizing MEGA’s computational efficiency is a promising avenue for future research [18].
Broader impact. MEGA contributes to the understanding of human perception from images. While there’s concern about potential misuse for intrusive surveillance, MEGA doesn’t reconstruct facial features, preserving anonymity. MEGA could have applications in healthcare, such as motor assessment of patients. This application would be positive, but potential prediction errors could negatively affect the care pathway. As with all deep learning models, training MEGA demands significant computing power, raising environmental concerns. Hyperparameter optimization wasn’t prioritized to minimize this cost, with model training taking about 9 days in total on 4 A100 GPUs for reported experiments.
Acknowledgments
This study is part of the EUR DIGISPORT project supported by the ANR within the framework of the PIA France 2030 (ANR-18-EURE-0022). This work was performed using HPC resources from the “Mésocentre” computing center of CentraleSupélec, École Normale Supérieure Paris-Saclay, and Université Paris-Saclay supported by CNRS and Région Île-de-France. We also thank Samir Sadok for his valuable help with the implementation of the masked autoencoder.
Appendix A Performance on an occlusion dataset
Table 3: Evaluation on 3DPW-OCC. We evaluate MEGA on an occlusion dataset and compare it to SOTA occluded HMR methods using standard metrics (see Sec. 4.1) in mm.
3DPW-OCC
Method Backbone PVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$
[…]
MEGA (ours) ResNet-50 93.8 79.8 51.5
MEGA (ours) HRNet-w48 78.9 66.3 43.7
[…]
Figure 6: Error distribution. We visualize the distribution of the error on 3DPW.
Appendix B Link between the deterministic and stochastic modes
@@
To gain deeper insights into the stochastic generation mode, we propose not only evaluating the best sample among the $Q$ generations (the common practice in the literature), but also assessing the mean of the generated meshes. In Tab. 4, we compare the performance of the average prediction for different $Q$ using the standard metrics (refer to Sec. 4.1). Additionally, we compute the Euclidean distance between the mean mesh and the one obtained in deterministic mode (Dist. to det., in mm) and the standard deviation of the predictions averaged over all the vertices (SD, in mm). To reduce the computational costs, this study is conducted on a randomly selected 10% subset of images from the 3DPW dataset. As $Q$ increases, the average prediction in stochastic mode approaches the deterministic prediction, with a distance around 7 mm for $Q=100$. The average prediction appears to converge toward a favorable solution, slightly outperforming the deterministic prediction in terms of PVE.
[…]
@@
When selecting the prediction with the minimum error among $N$ samples, we observe a shift in the distribution shapes, with errors concentrated toward lower values. While the highest error values decrease notably, the lowest remain relatively unchanged. This phenomenon likely occurs because the lowest errors typically correspond to meshes that are easier to predict and exhibit lower uncertainty. Consequently, the stochastic mode proves particularly beneficial for challenging images, where uncertainty is higher and multiple predictions offer valuable insights.
Appendix C Experimental details
[…]
@@ Appendix C Experimental details
HMR. For recovering human meshes from images in deterministic mode, we predict all images in a single step without randomness. In stochastic mode, we have to set the number of steps for generating the sequence of human mesh tokens and the amount of noise injected for the Gumble-max sampling. Note that we did not test MEGA with a ViT in stochastic mode. With HRNet, we generate meshes in 5 steps, and the initial noise temperature is 1. The generation process with ResNet-50 is made in 2 steps with an initial noise 10. The amount of noise at step $t$ is $A\times(1-\frac{t}{T})$ where $A$ is the initial noise temperature.
Figure 7: Prediction process iterations. We visualize the predictions for intermediate steps in stochastic mode (first 3 rows) and intermediate random generations (last 2 rows). All masked tokens are replaced by the first token of the codebook, corresponding to index 0.
@@
Random meshes generation. We generate random meshes in 20 steps. We want the generation to be completely random for the first steps so that the predictions are diverse. However, the last steps should be almost deterministic to obtain realistic meshes. The initial noise temperature is $A=1.2$, and the amount of noise at step $t$ is given by $(A\times(1-\frac{t}{T}))^{6}$.
Appendix D Generation process
We provide visualizations of the prediction process in stochastic mode and for random mesh generation. Specifically, for the HMR task, using the model with an HRNet backbone that generates meshes in 5 steps, given an image, we visualize all intermediate steps. In the case of random generation, we illustrate the meshes after 1, 5, 9, 13, 16, and 20 steps. At each step, all masked tokens are set to 0. Results are shown in Fig. 7. Importantly, our model does not predict unrealistic meshes in the first steps, as all masked tokens were replaced by the index 0.
For the HMR visualization, we observe that a rough estimate of the mesh is provided even in the initial steps, where only a few indices are predicted (3 during the first iteration and 7 during the second). This outcome is expected because the tokens set to be visible after each step are the most likely. Consequently, the indices with the highest certainty are selected first, enabling the construction of a preliminary mesh estimate. Subsequent steps refine these predictions, enhancing their realism.
@@ Appendix D Generation process
In random mesh generation, the initial mesh appears almost identical across all generations, as only one token is predicted in the first step (with all others set to 0 for visualization). However, diversity quickly emerges in subsequent steps, with the final steps refining the meshes to be more realistic. This pattern was anticipated, as the initial steps involve considerable randomness, whereas the later steps tend to become more deterministic.
Appendix E Qualitative results and failure cases
We present several failure cases in Fig. 8. Extreme poses can result in prediction errors, occasionally leading to non-anthropomorphic predictions due to the non-parametric nature of our approach. Notably, uncertainty is particularly high in such instances.
@@ Appendix E Qualitative results and failure cases
Figure 8: Failure cases. In failure cases, it is worth noting that our model predicts high uncertainty.
@@
Fig. 9 presents qualitative samples from in-the-wild datasets. We can observe that in some cases (for instance, images in the second row and left image in the fourth row), our predictions appear even more accurate than the ground truth. While this result is encouraging, it underscores the limited value of striving for fractions of millimeters of accuracy on datasets like 3DPW when the ground truth itself is imperfect. Additionally, we observe that uncertainty is significantly higher for occluded body parts, which often correspond to body extremities.
@@
Figure 9: Qualitative examples on 3DPW and EMDB. The first column is the groundtruth, the second and third are the image and the reprojection of the deterministic projection, and the third and fourth are uncertainty computed in stochastic mode and visualized in oriented and canonical points of view.
[…]

### v3

@@ 4.2 Deterministic inference mode
Quantitative evaluation. In  Tab. 1, we evaluate MEGA on the HMR task using the deterministic inference mode defined in Sec. 3.4, and compare its performance to SOTA methods on 3DPW [81] and EMDB [38]. Results for SOTA methods are taken from the corresponding papers when available or computed using official implementations and model weights. To ensure fairness, we only compare methods trained on standard datasets (see Sec. 4.1), with the exception of TokenHMR [22], which uses additional data but is included in the comparison because it also tokenized human pose. Note that the results for FastMETRO [13] and Virtual Marker [58] differ from those in their original papers because we present results without finetuning on the 3DPW dataset. Similar to [24, 22, 27], we chose not to use the 3DPW training set in order to measure generalization. Furthermore, it allows us to use the same model to evaluate MEGA on the occlusion dataset 3DPW-OCC [92] that contains sequences from the 3DPW training set.
MEGA outperforms all other methods on in-the-wild datasets. Using an HRNet [84] backbone, MEGA significantly surpasses both parametric [41, 49] and non-parametric [13, 58, 24] methods, especially on the EMDB [38] dataset. With a ViT [20] backbone, MEGA also achieves SOTA performance. However, comparisons with TokenHMR [22] are not completely fair, as this method is trained with additional 2D data and BEDLAM [7].
Figure 4: Qualitative samples. Given a single image with occlusions, MEGA makes diverse plausible predictions.
@@
MEGA det (ours) 81.6 68.5 44.1
Ablation study. Tab. 1 also provides the results of an ablation study. We tried a different distribution for the variable masking rate, such that $M=\lfloor N\tau\rfloor$ with $\tau$ uniformly sampled in [0,1[[0,1[[ 0 , 1 [ (“MEGA linear masking” in Tab. 1). Similar to [9], we find that the mask scheduling function improves the performance. This is also consistent with the findings of [31], which showed that a masking rate higher than 50% is beneficial for learning. We then tried masking 100% of the human mesh tokens during the supervised training phase (“MEGA full mask” in Tab. 1). We could expect this ablation to perform better as the training procedure corresponds to the deterministic inference mode. However, the results are slightly degraded, suggesting that sharing the same procedure between the self-supervised and supervised training stages is important. Additionally discarding the self-supervised pre-training of MEGA (“MEGA w/o pre-training + full mask” in Tab. 1) leads to a 2.5/6.0 mm increase in PVE on 3DPW/EMDB.
Results on an occlusion dataset. We quantitatively evaluate MEGA on the occlusion dataset 3DPW-OCC [92] in Tab. 2 to evaluate the robustness of our approach to occlusions. Despite not being tailored for HMR on occluded images, MEGA surpasses all other methods trained on the same data (see Sec. 4.1). This performance could stem from MEGA’s self-attention mechanism among mesh tokens. While visible parts can be predicted by leveraging image embeddings, occluded parts heavily rely on visible body parts for accurate inference.
@@ 4.3 Stochastic inference mode
Quantitative evaluation. We evaluate MEGA in stochastic inference mode (see Sec. 3.4) for the HMR task on the 3DPW [81] dataset, comparing our performance with SOTA multi-output approaches. Results from other methods are obtained from their respective papers. We exclusively compare methods trained on standard datasets (see Sec. 4.1), providing a version of MEGA using a ResNet-50 [30] backbone for fair comparisons. Following prior works [12, 6, 44], we assess accuracy using standard metrics computed with the minimum error sample out of $Q$ predictions, with $Q$ ranging from $1$ to $25$. We also calculate the relative improvement between 1 and 25 samples. Quantitative results are presented in Tab. 3 and qualitative samples are shown in Fig. 4.
MEGA achieves SOTA performance across all metrics and sample sizes. Notably, MEGA exhibits significantly higher accuracy than other methods when generating a single sample, and consistently demonstrates the best or second-best relative improvement among methods using a ResNet-50 backbone. With an HRNet [84] backbone, MEGA outperforms all probabilistic methods and surpasses SOTA single-output methods, even with a single stochastic generation. It is worth noting that it takes around 10 stochastic samples to outperform the deterministic generation (MEGA det), highlighting the utility of deterministic mode for quick and accurate predictions and the advantage of multiple predictions for enhanced accuracy. For a detailed comparison between deterministic and stochastic modes and a discussion on interpreting diverse predictions, please refer to Sec. A and Sec. D.
Generation process. We provide visualizations of the prediction process in stochastic mode. Specifically, using the model with an HRNet backbone that generates meshes in 5 steps (see Sec. B), given an image, we visualize all intermediate steps. Results are shown in Fig. 3. Importantly, our model does not predict unrealistic meshes in the first steps (all masked tokens were replaced by the index 0 for visualization purposes). We observe that a rough estimate of the mesh is provided even in the initial steps, where only a few indices are predicted (2 during the first iteration and 8 during the second as shown in Fig. 3). This outcome is expected because the tokens set to be visible after each step are the most likely. Consequently, the indices with the highest confidence are selected first, enabling the construction of a preliminary mesh estimate. Subsequent steps refine these predictions, enhancing their realism.
@@ 5 Conclusion
In this work, we explored self-supervised learning and masked generative modeling on human meshes for HMR. We introduced MEGA, a masked generative autoencoder designed to generate human meshes as discrete token sequences. MEGA’s flexible architecture and generation scheme enables the generation of diverse and realistic meshes and supports both single and multi-output HMR. An extensive evaluation demonstrates significant improvement over SOTA in both these domains.
Limitations. While MEGA generally produces accurate predictions, it struggles with extreme poses significantly divergent from the training data. Fig. 9 in Sec. E provides visualizations of these failure cases.
Future work. MEGA’s adaptable framework suggests potential applications beyond its current scope. Future research could explore generating human meshes conditioned on text inputs [18]. We could also complement image embedding with more observations, such as 2D pose tokens [26] or tokenized meshes of other individuals to model social interactions [29, 62]. Extending this work to videos by incorporating temporal masking during training [68, 69] or including more extreme poses in training data [77] may improve performance. Future works may also focus on other generative models such as discrete diffusion [3].
Broader impact. MEGA contributes to the understanding of human perception from images. While there is concern about potential misuse for intrusive surveillance, MEGA does not reconstruct facial features, preserving anonymity. MEGA could have applications in healthcare, such as motor assessment of patients. This application would be positive, but potential prediction errors could negatively affect the care pathway.
Supplementary Material
Figure 5: Error distribution. We visualize the distribution of the MPJPE in mm on the 3DPW dataset.
A Link between the deterministic and stochastic modes
@@
To gain deeper insights into the stochastic generation mode, we propose not only evaluating the best sample among the $Q$ generations (the common practice in the literature), but also assessing the mean of the generated meshes. In Tab. 4, we compare the performance of the average prediction for different $Q$ using the standard metrics (refer to Sec. 4.1). Additionally, we compute the Euclidean distance between the mean mesh and the one obtained in deterministic mode (Dist. to det., in mm) and the standard deviation of the predictions averaged over all the vertices (SD, in mm). To reduce the computational costs, this study is conducted on a randomly selected 10% subset of images from the 3DPW dataset. Tab. 4 shows that as $Q$ increases, the average prediction in stochastic mode approaches the deterministic prediction, with a distance around 7 mm for $Q=100$. The average prediction appears to converge toward a favorable solution, slightly outperforming the deterministic prediction in terms of PVE.
We also examine the distributions of the MPJPE on the 3DPW [81] dataset for MEGA in both deterministic and stochastic modes across various sample sizes $Q$. In stochastic mode, we analyze the average and best predictions. The results are reported in Fig. 5. Notably, the average prediction error of the stochastic mode appears to converge toward the deterministic prediction error, particularly as $Q$ increases; the distributions are very similar, with overlapping 95% confidence intervals. When $Q$ equals 1, the mean performance is comparatively lower, resulting in higher error values. These observations underscore the importance of having a deterministic mode for rapid and accurate predictions, which can be considered an estimator of the average prediction over $Q$ samples.
@@
When selecting the prediction with the minimum error among $N$ samples, we observe a shift in the distribution shapes, with errors concentrated toward lower values. While the highest error values decrease notably, the lowest remain relatively unchanged. This phenomenon likely occurs because the lowest errors typically correspond to meshes that are easier to predict and exhibit lower standard deviation. Consequently, the stochastic mode proves particularly beneficial for challenging images, where multiple predictions offer valuable insights.
B Experimental details
Pre-training stage. As detailed in the main body, the pre-training stage is done on a subset of AMASS [59], as introduced in [7]. We pre-train MEGA for 500 epochs, a task accomplished in less than a day on 4 A100 GPUs. MEGA is trained using the AdamW optimizer [57] with a cosine scheduler to adjust the learning rate. The base learning rate is $31e-$ 3, and we have a warmup of 20 epochs. The optimizer’s parameters are $\beta_{1}=0.9$, $\beta_{2}=0.99$, and the weight decay is 0.05.
Supervised training stage. The supervised training stage is done on a mix of standard datasets for HMR as presented in Sec. 4.1. We first train MEGA on MSCOCO [53] for 100 epochs and then train on the whole training set for 10 epochs. Each step takes about 1 day on 4 A100 GPUs. The training settings are exactly the same as the pre-training regarding learning rates and schedulers. For each training step, we start at epoch 0. Note that we have a lower learning rate in practice for the training on the mix of datasets because we stop the training before finishing the warm-up period. For training with HRNet and ResNet-50 backbones, the weights of the backbone are fine-tuned with the same settings as the other parameters of MEGA. When using ViT, the backbone is frozen during the training on MSCOCO, and we only fine-tune the last 10 blocks when training on standard datasets for computing power reasons.
@@ B Experimental details
HMR. For recovering human meshes from images in deterministic mode, we predict all images in a single step without randomness. In stochastic mode, we have to set the number of steps for generating the sequence of human mesh tokens and the amount of noise injected for the Gumble-max sampling. Note that we did not test MEGA with a ViT in stochastic mode. With HRNet, we generate meshes in 5 steps, and the initial noise temperature is 1. The generation process with ResNet-50 is made in 2 steps, and the initial noise temperature is fixed to 10. The amount of noise at step $t$ is $A\times(1-\frac{t}{T})$ where $A$ is the initial noise temperature [10, 9].
@@
Random meshes generation. We generate random meshes in 20 steps (see next section for more details). We want the generation to be completely random for the first steps so that the predictions are diverse. However, the last steps should be almost deterministic to obtain realistic meshes. The initial noise temperature is $A=1.2$, and the amount of noise at step $t$ is given by $(A\times(1-\frac{t}{T}))^{6}$.
C Random meshes generation
Figure 7: Prediction process iterations. We visualize the predictions for intermediate random generations. All masked tokens are replaced by the first token of the codebook, corresponding to index 0.
@@ C Random meshes generation
In Fig. 7, similar to the main paper, we visualize the generation process after 1, 5, 9, 13, 16, and 20 steps. The initial mesh appears almost identical across all generations, as only one token is predicted in the first step (with all others set to 0 for visualization). However, diversity quickly emerges in subsequent steps, with the final steps refining the meshes to be more realistic. This pattern was anticipated, as the initial steps involve considerable randomness, whereas the later steps tend to become more deterministic.
Figure 8: Visualization of the predictions diversity. We visualize the standard deviation of the 3D location of each vertex. Bluish regions in the mesh indicate low standard deviation, while reddish areas signify higher standard deviation.
D Interpreting diversity in predictions
In the stochastic mode, MEGA makes diverse predictions given a single image. After making multiple predictions given an image, we can compute the standard deviation of the position of each vertex and interpret this value as a measure of the uncertainty in predictions. Indeed, if all samples are similar, we can conclude that the model is “certain” about this prediction. When the results are very diverse given a single image, we can interpret that as high uncertainty.
Some qualitative samples are shown in Fig. 8. The two images on the top show non-occluded bodies, and the relative depth between body parts is easy to perceive. Thus, the standard deviation is low; the model consistently makes accurate predictions. The body is partially occluded in the two bottom images, and the depth of some body parts (such as the left arm in the left image) is hard to estimate. The model makes diverse predictions, which can be interpreted as high uncertainty.
@@ D Interpreting diversity in predictions
Figure 9: Failure cases. In failure cases, it is worth noting that our model predicts very diverse results, which can be interpreted as high uncertainty.
E Qualitative results and failure cases
We present several failure cases in Fig. 9. Extreme poses can result in prediction errors, occasionally leading to non-anthropomorphic predictions as we do not rely on the SMPL parameters [56]. Notably, the standard deviation of the vertices’ position is exceptionally high in such instances.
@@ E Qualitative results and failure cases
Fig. 10 presents qualitative samples from in-the-wild datasets. We can observe that in some cases (for instance, images in the third and fourth rows on the right), our predictions appear even more accurate than the ground truth. While this result is encouraging, it underscores the limited value of striving for fractions of millimeters of accuracy on datasets like 3DPW when the ground truth itself is imperfect.
@@
Figure 10: Qualitative examples on 3DPW and EMDB. The first column is the groundtruth, the second and third are the image and the reprojection of the deterministic prediction, and the fourth is the prediction using the deterministic mode of MEGA with an HRNet backbone.